OpenStudy (anonymous):

((3)/(2v))+((3v)/(4v^2-2v))

OpenStudy (astrophysics):

\[\frac{ 3 }{ 2v } + \frac{ 3v }{ 4v^2-2v }\] and what do you want to do with this?

OpenStudy (anonymous):

simplify

OpenStudy (astrophysics):

Use the following algebra property \[\frac{ a }{ b } \pm \frac{ c }{ d } = \frac{ ad \pm bc }{ bd }\]